The perimeter of a regulation singles tennis couurt is 210ft. and the length is 30ft. less than 4 times the width. What are the dimensions of the court?
Answer by josgarithmetic(39629)

the length is 30ft. less than 4 times the width.
'
the x is 30ft. less than 4 times the y.
x=-30+4y
x=4y-30

Length is x and width is y

2x+2y=210
2(4y-30)+2y=210
8y-60+2y=210
10y-60=210
10y=210-60
y=21-6
y=15 and x=4*15-30=30
Meaning length is 30 ft, and width is 15 ft.